OpenStudy (anonymous):

ERROR ANALYSIS: Below is a bonus question that appeared on the algebra test and a student's response. The ratio of junior varsity members to varsity members on the track team is 3:5. There are 24 members on the team. Write a proportion to find the number of junior varsity members. 3/5 = x/24 The student did not receive the bonus points. Why is this proportion incorrect?

OpenStudy (anonymous):

he did not actually solve for the number of juniors

OpenStudy (anonymous):

Should it be 3/8 = x/24

OpenStudy (anonymous):

No the proportion is incorrect it doesn't say to solve it.

OpenStudy (anonymous):

3 out of 5 not 3 out of 8

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

But then why is 3/5 = x/24 incorrect.

OpenStudy (anonymous):

going for coffee, nvm, disregard

OpenStudy (anonymous):

3/8 because if there is 3 juniors and 5 varsity then the total is 8.

OpenStudy (phi):

There are two types junior varsity and varsity together they add up to 24 members they give you the ratio of junior : varsity = 3:5 if you want the ratio of junior to total, you write 3: 8 (where 8= 3+5) now you can say x/24= ⅜

OpenStudy (anonymous):

phi so it is 3/8 =x/24

OpenStudy (phi):

yes, you got it correct.